OpenStudy (anonymous):

find the y-intercept of this equation 4x-y=0

OpenStudy (anonymous):

show your work

OpenStudy (anonymous):

Put x=0, and you get the y-intercept. \[4*0 - y = 0\] Therefore your y-intercept is \[y=0\] Your x-intercept is zero as well because this equation only touches the axes at zero.

OpenStudy (anonymous):

straight line going up right

OpenStudy (anonymous):

Yeah, anytime you see an equation that has only an y and an x in it, like 3x+4y=10 - it is always a line. If x is in the power of something like \[3x^2+4y=10\]thats where math gets a bit interesting.